OpenStudy (anonymous):

13x+24=-41 use the distributive propery to simplify or completely solve PLZ HELP!!!!!!!

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

-5!

OpenStudy (anonymous):

So sorry

OpenStudy (anonymous):

that was fast

OpenStudy (anonymous):

This is simple stuff. I'm on stuff like: If the complex number x = 3 + bi and |x|2 = 13, which is a possible value of b?

OpenStudy (anonymous):

Which I'm currently trying to get help with.

OpenStudy (anonymous):

thats confusing

OpenStudy (phi):

add -24 to both sides, like this: 13x+24 + -24=-41+ -24 can you simplify both sides ?

OpenStudy (anonymous):

we figured it out